Two angles are complementary. The measure of one angle is 21 more than twice the measure of the other. Find the angles
Answer by JulietG(1812)

Complementary angles add up to 90 degrees
A + B = 90
A = 2B+21
Substitute the value of A from the second equation into the first
A + B = 90
(2B+21) + B = 90
3B + 21 = 90
3B = 69
B = 23
If B is 23, then A is (2*23)+21, or 46+21, or 67